# Collaborative Document Editing with Multiple Users and AI Agents

### v1

@@ 6.3.4. Final Texts
The final texts had a mean length of 1623 words (SD=1306.93, min=226, max=4767). To assess the readability of texts, we computed the Flesch reading-ease score (Flesch, 1948). We removed four texts for this analysis since they included non-English parts. The mean readability score of 17.71 (SD=13.71, min=2.87, max=38.65) can be interpreted as “very difficult to read” and is described as appearing in “Scientific” and ”Academic” reading (Flesch, 1948), reflecting the nature of most of the groups’ writing projects.
@@ 6.4. Interviews and Open Comments
We transcribed the interviews with Microsoft Teams and analysed them following open and axial coding principles (Corbin and Strauss, 2014), verifying transcripts against audio where needed. One author iteratively grouped, merged, and split codes into categories and subcategories. Another author reviewed and refined the outcome, also going back to the recordings where needed and involving the team for reaching consensus. We coded open comments from the final questionnaire in the same way. Below, we present the main themes we identified.
@@ 6.4.1. Territorial functioning: An agent belongs to its creator
Agents and tasks were seen as belonging to their respective creators, with all groups except for two expressing this clearly in the interviews. Their statements reveal different assumptions and expectations that explain this.
One rationale is that everyone knows best what they want from AI support; as $P_{2,4}$ said “we had two or three different agents, so I had one more which I wanted to focus more on […] fixing typos or grammar.” And $P_{1,5}$ found that “you cannot really decide how many AI assistants someone would need”. Similarly, $P_{2,2}$ explained: “We have different AI agents and we can set different tools and […] we can use them for a different purpose.” $P_{1,7}$ valued this for prompting: “I know what to expect. It’s better to use an agent that I made myself, that I know […] So I also then know how, or I have a better feeling over how, I write the prompts in the comments.” $P_{2,13}$ even envisioned individual visibility: “Something that might be useful for one author might actually be a hindrance to another author, so […] people could have their own agents that only show up in their version of the document or on their screen.”
@@
Finally, workflows were another reason for creating agents individually. As $P_{2,11}$ said: “I have certain ways of organising myself […] so therefore I think it would be [better] for my workflow.” Group 12 expressed a similar view, group 4 decided upfront to split the work, group 13 worked sequentially on their task, and group 1 described they explored the new features individually.
@@ 6.4.2. Emerging roles and norms: Sharing agents is caring
Moreover, people created agents and tasks with collaboration in mind. For example, $P_{1,8}$ was keen to “make sure that the agent in the end is reliable and it’s a good team player for both of us.” And $P_{2,2}$ described it for tasks: “We didn’t split it one task per person. We just kind of mix it.” $P_{2,11}$ saw individually owned agents as an optional team offer: “I’d be happy to use shared ones, but […] if I don’t like them, then I still want to have my own. […] I’d be happy to share my stuff and people should only use it if they find a benefit.” And $P_{1,7}$ reflected: “As a team, you could have this agent. Everybody knows what it’s for and you know you use it for this task. So of course, not everyone has to make their own five agents.”
A few people used agents/tasks as a mediator or “glue” in the team. For example, $P_{2,4}$ used the all-users-offline task trigger to get recaps: “After logging in again it created the summary which was quite nice. So just to get a recap of what had happened so far.” $P_{2,12}$ reflected on interaction with agents as available alternatives for delegating tasks: “I find it really useful to [use tasks and comments for agents] because I think it’s a natural way when you work together […] Yeah, you write a task for another co-author […] and so I think I made no difference if I assign a task to [him] or to an AI agent. I think it was, from the feeling, nearly the same.” Related, $P_{1,7}$ liked how they could use an agent to keep work coherent that they had split across sections with their partner: “I’ve read what he’s done and then I could come back and then ask [the] agent: He wrote something like this. Can you put something related now in my text?”
@@ 6.4.3. One vs multiple agents: A tradeoff between expected functional and perspective values
We observed a specific strategic tradeoff: People weighed the setup effort against the expected benefits of multiple agents. Different views were discussed within nine of the teams. Overall, 14 participants brought up rationales in favour of one agent, and 20 (also) in favour of multiple ones.
@@
In contrast, those inclined to use multiple agents focused on perspective value, such as $P_{1,12}$: “I also really liked using AI agents to incorporate viewpoints that are not my own.” Similarly, $P_{1,1}$ said: “You can use the second agent to check the text, like, to play as a professor, as a reviewer.” Group 12 also liked agents as personas for feedback. $P_{2,4}$ liked “having different agents for different things”, also echoed by $P_{2,13}$ and $P_{2,14}$. $P_{2,8}$ compared it to human teams: “When you are collaborating with others in the same team, you know, some people are more good at this. Some people are more good at that. So for agents it’s the same.” $P_{2,10}$ reflected on both value and effort: “I can have an external reviewer that reviews from the lens of that community. […] I’m a bit sceptical [about] having to create and manage and stuff.”
@@
A related rationale was that agent roles support prompting. For instance, $P_{2,7}$ said that “it could be faster then to refer to specific agents […] and not explain the only one AI agent [everything]. It’s like working in a company. Everybody has it’s role.” Similarly, $P_{1,3}$ considered that defined roles shorten task prompts: “It can be beneficial to have several agents predefined because their answers can be different. Sometimes I want to create a very, very specific description. I don’t want to write a very long task for that.” $P_{1,7}$ also found that “defining the agent spares you a part of the prompts writing.” And $P_{1,13}$ said: “I first thought of the persona pattern [for prompts] and I would try to push the LLMs as much as possible into this role. […] I thought of tasks I would assign to the agents and design the agents the way I want to implement the tasks.”
In summary, participants weighed setup effort against benefits of multiple agents, with their conclusion depending on their focus on functional vs perspective values and future prompting.
6.4.4. Feedback on the design features
Participants shared rich feedback on the design features and ideas for further features. For brevity, here we focus on their key pros and cons.
Participants liked that direct editor integration saves time and effort for interacting with AI. For example, $P_{2,1}$ found that “it saves a lot of time that I don’t have to switch tabs.” $P_{1,2}$ had “a great experience to use it directly in the document, and it saves a lot of time [and] it was surprisingly good.” And $P_{1,5}$ said the “assistants were pretty useful […] Previously, I had to divide my text into several fragments if I wanted to check it in ChatGPT.” $P_{2,3}$, $P_{3,4}$, $P_{1,9}$, and $P_{2,12}$ made similar direct comparisons to their usual AI tools.
@@ 6.4.4. Feedback on the design features
The comment integration of agents was valued here specifically. As $P_{2,8}$ said: “I really like the comments. You can add the agent directly in the comments and that was very convenient, and it feels like you’re actually talking to a person or like you’re discussing with someone.” Similary, $P_{2,3}$ found the comments “very helpful, especially to collaborate with others […] and very integrated.” $P_{1,7}$ wanted to keep using this feature: “What I would include in the usual editor is the comments with the AI agents.” The comments integrating AI were also highlighted as a valued feature by $P_{1,1}$, $P_{1,3}$, $P_{2,9}$ and everyone in groups 4, 8, 11, 12, and 13.
@@
For some, using the profile UI was easier than figuring out what information to enter. As $P_{1,6}$ said, “AI agent creation [is] easy to understand” and $P_{1,11}$ found it “really easy to get your head around, to imagine that you have this little team”, while $P_{2,4}$ “found it a bit confusing to decide whether it now was a skill or just a background.” $P_{1,3}$ reflected explicitly on both sides: “I did not struggle to create agents. Basically, because the interface for agent creation was really easy to understand, but like, some intuition about what kind of agent I want to create. That’s actually kind of [the] problem.” $P_{1,13}$ generalised it: “It’s the typical issue with LLMs. If you’re not really sure what you want to generate, then the answers are typically not so high quality.”
Overall, participants preferred to give tasks via manual triggers and comments. One reason was the lack of progress feedback or the inherent invisibility of automated triggers such as leaving the document. For instance, $P_{2,5}$ said: “It works, but I have seen no results, but I have not understood why, but only yesterday, I see it was there.” Another reason was that calling agents manually in comments constrains their response to one new comment in the expected location. For example, $P_{1,13}$, $P_{2,4}$, and $P_{1,3}$ highlighted how it was useful to “tag an agent and ask him a question regarding the selected text” ($P_{1,3}$). Related, $P_{2,13}$ liked this for task shortcuts: “I really like the feature where you […] select the text […] and just use a task you defined beforehand.”
Some reflected more explicitly on trust and control for preferring manual triggers, such as $P_{1,14}$ (“You don’t have much control of what’s going on.”) and $P_{1,7}$: “You’re always afraid that some changes will happen to your documents. […] So I don’t think people will be doing a lot of really autonomous stuff.” $P_{2,10}$’s comment echoes this: “I only did manual because I really hated the idea of it doing stuff when I was not there. […] I just see it too much as, like, I merge and I control it. I didn’t want it to do anything autonomously.” In contrast to these views, only one person ($P_{1,12}$) envisioned giving a lot more initiative to agents, to “feel like the AI is my colleague”.
@@
Finally, the main point of critique concerned the LLM’s output. Group ten was the most critical; it did not meet the bar for them, as they perceived it as worse than ChatGPT. All other groups found it useful overall but voiced the specific concern that AI output was too verbose, in three flavours: First, AI agents added too many comments, when not tagged in a specific place via a comment but rather given a task via the task list. As $P_{1,3}$ concluded “it’s left, like, comments over all [the] document and I would not want to use these autonomous tasks too often.” Second, some AI comments were longer than desired, which made $P_{1,4}$, for example, suggest to “shorten the comments quite a bit”. Third, AI responses included unwanted conversational parts (e.g. “Sure, here’s a draft: …”), such as in the case of $P_{1,6}$: “[agents] reply in conversational messages, [which would be] added to the text, regardless of whether you choose to append or replace.”
In summary, participants valued the integrated, collaborative AI features, while pointing to challenges in agent definition and control, as well as tuning the volume of AI contributions.
@@ 6.5. Further Text Analyses
We further analysed the texts entered in the agent profiles, task descriptions, and the comments that delegated ad-hoc tasks to agents. We deductively coded co-creative roles for the agent profiles, and the user’s primary intent and intended writing stage for task descriptions and delegations in comments. Possible codes were the typical writing stages (outlining, drafting, revising), aligned with the literature (Lee et al., 2024). Similarly, we defined co-creative roles based on related work (Lowry et al., 2004; Posner and Baecker, 1992), as well as the primary intents (Yuan et al., 2022; Dang et al., 2023; Gero et al., 2022) (see below for concrete codes). Three researchers coded the data logged from the first six groups and then discussed the codes to find consensus and to refine the codebook. Each coder finalized the coding of the dataset independently. Finally, the lead coder streamlined the codes.

### v2

@@ 6.1.6. Final texts
The final texts had a mean length of 1623 words (SD=1306.93, min=226, max=4767). To assess the readability of texts, we computed the Flesch reading-ease score (Flesch, 1948). We removed four texts for this analysis since they included non-English parts. The mean readability score of 17.71 (SD=13.71, min=2.87, max=38.65) can be interpreted as “very difficult to read”. For contextualisation, these scores fit to those reported in analyses of contemporary academic literature (Plavén-Sigray et al., 2017; Yeung et al., 2018) and are described as appearing in “Scientific” and ”Academic” texts (Flesch, 1948), reflecting the nature of most of the groups’ writing projects.
@@ 6.2. Interviews and open comments
We transcribed the interviews with Microsoft Teams and analysed them following open and axial coding principles (Corbin and Strauss, 2014), verifying transcripts against audio. One author iteratively created, grouped, merged, and split codes into (sub)categories. Another reviewed and refined this, going back to the recordings where needed and involving the team for consensus. We similarly coded open comments from the final questionnaire.
@@ 6.2.1. Territorial functioning: An agent belongs to its creator
Agents and tasks were seen as belonging to their respective creators, with all groups except for two expressing this. One rationale for this is that everyone knows best what they want from AI support; as $P_{2,4}$ said “we had two or three different agents, so I had one more which I wanted to focus more on […] fixing typos or grammar.” And $P_{1,5}$ found that “you cannot really decide how many AI assistants someone would need”. Similarly, $P_{2,2}$ explained: “We have different AI agents and we can set different tools and […] we can use them for a different purpose.” $P_{1,7}$ valued this for prompting: “I know what to expect. It’s better to use an agent that I made myself, that I know […] So I also then know how, or I have a better feeling over how, I write the prompts in the comments.” $P_{2,13}$ even envisioned individual visibility: “Something that might be useful for one author might actually be a hindrance to another author, so […] people could have their own agents that only show up in their version of the document or on their screen.”
@@
Finally, workflows that teams approached the study with were another reason. As $P_{2,11}$ said: “I have certain ways of organising myself […] so therefore I think it would be [better] for my workflow.” Group 12 expressed a similar view and group 1 described they explored the new features individually.
@@ 6.2.2. Emerging roles and norms: Sharing agents is caring
People created agents and tasks with collaboration in mind. For example, $P_{1,8}$ was keen to “make sure that the agent in the end is reliable and it’s a good team player for both of us.” And $P_{2,2}$ described it for tasks: “We didn’t split it one task per person. We just kind of mix it.” $P_{2,11}$ saw individually owned agents as an optional team offer: “I’d be happy to use shared ones, but […] if I don’t like them, then I still want to have my own. […] I’d be happy to share my stuff and people should only use it if they find a benefit.” And $P_{1,7}$ reflected: “As a team, you could have this agent. Everybody knows what it’s for and you know you use it for this task. So of course, not everyone has to make their own five agents.”
@@ 6.2.3. One vs multiple agents: A tradeoff between expected functional and perspective values
We observed a specific strategic tradeoff: People weighed the setup effort against the expected benefits of multiple agents. Different views were discussed within nine of the teams. Overall, 14 participants brought up rationales in favour of one agent, and 20 (also) in favour of multiple ones. We report on their rationales in detail next:
Rationales for one agent
@@
Those inclined to use multiple agents focused on perspective value, such as $P_{1,12}$: “I also really liked using AI agents to incorporate viewpoints that are not my own.” Similarly, $P_{1,1}$ said: “You can use the second agent to check the text, like, to play as a professor, as a reviewer.” Group 12 also liked agents as personas for feedback. $P_{2,4}$ liked “having different agents for different things”, echoed by $P_{2,13}$ and $P_{2,14}$. $P_{2,8}$ compared it to human teams: “When you are collaborating with others in the same team, you know, some people are more good at this. Some people are more good at that. So for agents it’s the same.” $P_{2,10}$ reflected on both value and effort: “I can have an external reviewer that reviews from the lens of that community. […] I’m a bit sceptical [about] having to create and manage and stuff.”
@@
A related rationale was that agent roles support prompting: $P_{2,7}$ said that “it could be faster then to refer to specific agents […] and not explain the only one AI agent [everything]. It’s like working in a company. Everybody has its role.” Similarly, $P_{1,3}$ considered that defined roles shorten task prompts: “It can be beneficial to have several agents predefined because their answers can be different. Sometimes I want to create a very, very specific description. I don’t want to write a very long task for that.” $P_{1,7}$ also found that “defining the agent spares you a part of the prompts writing.” And $P_{1,13}$ said: “I first thought of the persona pattern [for prompts] and I would try to push the LLMs as much as possible into this role. […] I thought of tasks I would assign to the agents and design the agents the way I want to implement the tasks.”
Summary
Participants weighed setup effort against benefits of multiple agents. What drives their varying preferences is their individual focus on either functional or perspective values and future prompting. Notably, people did not mention specific writing use cases in their rationales (e.g. one agent for emails, while a paper needs more), suggesting that they made these considerations on a more general level of what they expect from AI in writing.
6.2.4. Collaborative processes with AI agents
Eight groups worked synchronously, eight (also) asynchronously. The groups also used chat apps and video calls. Six groups split the writing by “location” (e.g. one section per person). Thus, our sample covers using agents, tasks, and comments in various collaborative processes.
AI comments as team “action items”
Consistently throughout the groups, AI comments became shared material (Section 6.2.2), as participants applied or rejected AI suggestions in comments, regardless of who had created or triggered the agent (Section 6.1.5). Thus, these decisions on AI suggestions became a new, shared “action item” in the collaborative process. This was an overarching aspect, not tied to a specific moment or stage of writing.
Using tasks for planning and structuring
The shared task list (Section 6.2.2) was integrated into collaborative processes in a forward-thinking way: For example, G1 created tasks for planning, while working synchronously. In contrast, G11 worked asynchronously and in the interview reflected on using tasks for structuring a document first (“For this overall structure […], top down breaking things down […] I think tasks would work better.”, $P_{2,11}$). And for G12, who worked asynchronously and also in parallel, the tasks for agents felt similar to writing todos to each other. Thus, the task list feature was integrated into team processes with a forward-thinking perspective, in contrast to the more ad-hoc comments.
Agents as ”glue” for collaborative processes
A few people supported asynchronous collaboration with agents and tasks. For example, $P_{2,4}$ used the all-users-offline task trigger to see recaps via AI comments when coming back: “After logging in again it created the summary which was quite nice. So just to get a recap of what had happened so far.” $P_{2,12}$, also working asynchronously, reflected on interaction with agents as available alternatives for delegating tasks: “I find it really useful […] because I think it’s a natural way when you work together […] Yeah, you write a task for another co-author […] and so I think I made no difference if I assign a task to [him] or to an AI agent. I think it was, from the feeling, nearly the same.” $P_{1,7}$ used an agent to keep work coherent that they had split across sections with their partner: “I’ve read what he’s done and then I could come back and then ask [the] agent: He wrote something like this. Can you put something related now in my text?” And in G5, who split the work into sections to work on asynchronously, $P_{1,5}$ created a “praiser” agent for motivation in the absence of their partner. In summary, in such cases, the agents were used to support “handover” or “partner absent” moments of asynchronous workflows.
Using (custom) agents in roles specific to writing stages
Many considered agents to address specific stages in their collaborative workflow. For example, G2 used them to create a first draft synchronously, while G3 and G4 asynchronously created custom “reviewer” agents, as also reflected on in G10 (“I can have an external reviewer that reviews from the lens of that community”, $P_{2,10}$). G8 designated one user to setup all agent roles for the team. In such cases, (custom) agents were brought in with awareness of the team’s (negotiated) workflows and with writing process stages in mind.
Concerns about workflows with agents
Some reflected on downsides for collaboration: G10 commented that agents cannot take responsibility for mistakes. $P_{2,8}$ felt they interacted more with agents than their partner at times, despite sharing a video call (“It feels like I’m still more, like, interacting with the AI agents, rather than the people”). $P_{2,13}$ expected conflicts when one user deactivates an auto-triggered task that another finds helpful. And G4 used the “all users offline” trigger to avoid getting interrupted by agent comments with other auto-triggers. In summary, the concerns reveal reflections on possible distractions and interruptions, as well as questions of responsibility.
6.2.5. Values and demands of the design features
Participants’ rich feedback, combined with logging data and ratings (Section 6.1.1, 6.1), enables us to dissect the values and demands of the features.
Value: Integration compared to known designs
Participants valued the direct editor integration. Their comparisons to their usual work highlight two aspects: One valued aspect is avoiding the need for switching between the writing environment and AI, including copy-paste: $P_{1,5}$ found the “assistants were pretty useful […] Previously, I had to divide my text into several fragments if I wanted to check it in ChatGPT.” And $P_{3,4}$ found it “more comfortable than copying the stuff over to, for example, ChatGPT.” $P_{1,9}$ also shared this value: “You can use AI agents to brainstorm new ideas. You don’t have to go to GPT.” A related aspect is that this integration is perceived as fast, in comparison to separate AI tools: As $P_{2,3}$ said: “I could do everything much faster without using ChatGPT or other applications […] because I could use these agents.” And $P_{1,2}$ had “a great experience to use it directly in the document, and it saves a lot of time [and] it was surprisingly good.” $P_{2,7}$ shared a similar view: “What I like about it, it was helping to make it really faster than I will do it without it.” And $P_{2,1}$ said “it saves a lot of time that I don’t have to switch tabs. I often use […] other online tools. And in this app […] I could just ask the agent.” $P_{2,12}$ reflected on both aspects: “Previously my workflow was that I take my text, copy it to ChatGPT or other chatbots, and write, each time again, my prompt or copy the prompt, and ask GPT to do that. And with the agent it’s much easier and faster because I can […] work with different agents in parallel. That I found really, really helpful.”
Value: Using comments with agents
@@ 6.2.5. Values and demands of the design features
The design integrating agents specifically into comments was valued, as explicitly highlighted by $P_{1,1}$, $P_{1,3}$, $P_{2,9}$ and everyone in groups 4, 8, 11, 12, and 13. For example, $P_{2,8}$ said: “I really like the comments. You can add the agent directly in the comments and that was very convenient, and it feels like you’re actually talking to a person or like you’re discussing with someone.” Similary, $P_{2,3}$ found the comments “very helpful, especially to collaborate with others […] and very integrated.” $P_{2,11}$ valued the AI comments as additions in comparison to Google Docs: “I like the overall […] similarity to Google Docs style because, I mean, for me it was very easy to to get into it […] and then having AI integrated in the form of […] comments and to have these tasks. […] I really like to have AI integrated in the tool.” $P_{1,7}$ wanted to keep this feature when reflecting on their usual tools: “What I would include in the usual editor is the comments with the AI agents.”
Demand: Setup of agent profiles
@@
For some, using the profile UI was easier than figuring out what information to enter. As $P_{1,6}$ said, “AI agent creation [is] easy to understand” and $P_{1,11}$ found it “really easy to get your head around, to imagine that you have this little team”, while $P_{2,4}$ “found it a bit confusing to decide whether it now was a skill or just a background.” $P_{1,3}$ reflected explicitly on both sides: “I did not struggle to create agents. Basically, because the interface for agent creation was really easy to understand, but like, some intuition about what kind of agent I want to create. That’s actually kind of [the] problem.” $P_{1,13}$ generalised it: “It’s the typical issue with LLMs. If you’re not really sure what you want to generate, then the answers are typically not so high quality.” In summary, agent setup was a metacognitive demand (cf. (Tankelevitch et al., 2024)) rather than a usability demand, fitting to prompting challenges in related work (Benharrak et al., 2024; Zamfirescu-Pereira et al., 2023).
Demand: Verbose AI output
@@
The main point of critique concerned LLM output. G10 was the most critical; it did not meet the bar for them, as they perceived it as worse than ChatGPT. All other groups found it useful overall but voiced the concern that AI output was too verbose, in three flavours: First, AI agents added too many comments, when not tagged in a specific place via a comment but rather given a task via the task list. As $P_{1,3}$ concluded “it’s left, like, comments over all [the] document and I would not want to use these autonomous tasks too often.” Second, some AI comments were longer than desired, which made $P_{1,4}$, for example, suggest to “shorten the comments quite a bit”. Third, AI responses included unwanted conversational parts (e.g. “Sure, here’s a draft: …”), such as in the case of $P_{1,6}$: “[agents] reply in conversational messages, [which would be] added to the text.”
Addressing the demands and control needs with manual triggers
Overall, participants preferred to give tasks via manual triggers and comments. One reason was that calling agents manually in comments constrains their response to one new comment in the expected location. This also addresses AI verbosity reported above via targeted requests. For example, $P_{1,13}$, $P_{2,4}$, and $P_{1,3}$ highlighted how it was useful to “tag an agent and ask him a question regarding the selected text” ($P_{1,3}$). Related, $P_{2,13}$ liked this for task shortcuts: “I really like the feature where you […] select the text […] and just use a task you defined beforehand.” Another reason was the lack of progress feedback or the inherent invisibility of automated triggers such as leaving the document. For instance, $P_{2,5}$ said: “It works, but I have seen no results, but I have not understood why, but only yesterday, I see it was there.”
Some reflected on trust and control for preferring manual triggers, such as $P_{1,14}$ (“You don’t have much control of what’s going on.”) and $P_{1,7}$: “You’re always afraid that some changes will happen to your documents. […] So I don’t think people will be doing a lot of really autonomous stuff.” $P_{2,10}$’s comment echoes this: “I only did manual because I really hated the idea of it doing stuff when I was not there. […] I just see it too much as, like, I merge and I control it. I didn’t want it to do anything autonomously.” In contrast to these views, only two participants ($P_{1,12}$, $P_{1,4}$) envisioned giving more initiative to agents, for example to “feel like the AI is my colleague” ($P_{1,12}$).
Summary
We identified specific values and demands of the design features: (1) Integrating AI agents into collaborative documents is valued for avoiding frequent app switches and copy-pasting back-and-forth. (2) Agent profiles are perceived as easy to use, specifically the structured CV, yet impose metacognitive demands (knowing what you need). (3) The comment feature is valued as the specific way of integrating agents, yet a high number or verbose comments are demanding. (4) The feature for prompting agents ad-hoc by selecting text and writing a comment is valued, and so is the feature for triggering a defined task manually (from either task list or created toolbar shortcut button). These help address the demands by constraining AI output to one expected place (selected text). This is preferred over the automated trigger feature, also for reasons of trust and control.
@@ 6.3. Further Text Analyses
We further analysed the texts entered in agent profiles, task descriptions, and comments that delegated ad-hoc tasks to agents. We deductively coded co-creative roles for the agent profiles, and the user’s primary intent and intended writing stage for task descriptions and delegation comments. Possible codes were the typical writing stages (outlining, drafting, revising) (Lee et al., 2024). Similarly, we defined co-creative roles based on related work (Lowry et al., 2004; Posner and Baecker, 1992), as well as the primary intents (Yuan et al., 2022; Dang et al., 2023; Gero et al., 2022) (see below for concrete codes). Three researchers coded the data logged from the first six groups and then discussed the codes to find consensus and to refine the codebook. Each coder finalised the coding independently. Finally, the lead coder streamlined the codes.